• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name K. Kumar

    Log entry time 23:16:05 on August 21, 2009

    Entry number 285180

    keyword=Swing shift summary

    16:00 At start of shift, we are trying to find out if the mapper if functional

    16:30 Kent, Luis and Mark go in to reboot the quads.

    17:00 beam is down due to a vacuum glitch

    18:00 Request for controlled access, since beam is still down. Mapper is still not functional, but we do not have the personnel to work on on it right now, so we punted on this.

    18:10 Ramesh and Dustin enter the hall to fix the right arm trigger

    19:45 Ramesh, Dustin and Alex are done in the right arm, but beam is at least an hour away, so they are fine tuning things in the left arm.

    20:00 Beam to be available no later than 11 pm, but likely not before 10 pm

    20:25 Ramesh and Dustin are done setting up the trigger but Ramesh has to go back in to plug in ribbon cables to scalars in the left arm

    21:15 Ramesh is done

    22:20 Alex will take a brief controlled access to power cycle the compton laser power supply

    22:30 Alex is done and we are in beam permit awaiting beam